OpenStudy (anonymous):

I need to find x in this equation here: 47.81 = 63.466e^0.61119x ....how??? Please actually explain what's going on rather than just giving me the answer too, I'd like to understand what's going on. Thanks!

OpenStudy (anonymous):

Oops, typo. The exponent should be 0.6119x. Not 0.61119x.

OpenStudy (anonymous):

I think it should be: ln 47.81 = ln 63.466 +0.6119x (ln 47.81-ln 63.466)/0.6119 = x (ln (47.81/63.466))/0.6119 = x